You need P5,000 per year for four years to go to college, your father invested P6,000 in 7% account for your education when you were born. If you withdraw at the end of your 17th, 18th, 19th and 20th birthday, how much will be left in your account at the end of your 21st birthday.

1
Expert's answer
2022-03-07T12:11:16-0500

FI"=P1(1+I)^n"

"=6000(1+0.07)^{21}"

"=2484.34"

"F2=\frac{A(1+I)^n-1)}{0.07}"

"F2=\frac{5000[(1+0.07)^{4}-1]}{0.07}" =1553.98

"F3=F2(1+I)" =1553.98(1+0.07)=1662.75

Money left==F1-F3=2484.34-1662.75=

821.2

So he will be left with P 821.2